OpenStudy (anonymous):

How would you solve the intigral from -1 to 2 for the equation (x/(x^2+1))dx?

OpenStudy (bahrom7893):

let u =x^2 +1 du = 2xdx rewrite the integral as: 1/2 Integral (2xdx/(x^2+1))

OpenStudy (bahrom7893):

1/2 Integral (du/u) = 1/2 Ln|u| = 1/2 Ln(x^2+1) + C

OpenStudy (bahrom7893):

Note you dont need abs value any more because x^2 is always positive and positive plus one is still positive!

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

TYVM

OpenStudy (anonymous):

wait... is the anser 1/2ln5/2 or 1/2ln3?

OpenStudy (bahrom7893):

oh sorry forgot to evaluate: 1/2 Ln(x^2+1) from -1 to 2: (1/2){Ln(2+1) - Ln(1+1)} = (1/2) {Ln3 - Ln2} = (1/2)Ln(3/2)

OpenStudy (bahrom7893):

oh wait no

OpenStudy (bahrom7893):

1/2ln5/2

OpenStudy (bahrom7893):

2^2 is 4, not 2... (1/2){Ln(4+1) - Ln(1+1)} = (1/2) {Ln5 - Ln2} = (1/2)Ln(5/2)

OpenStudy (anonymous):

yep ok that helps alot ty